OpenStudy (anonymous):

find the area of the surface generated by revolving the curve about the indicated axis. y=(x^3)/14 x=0 x=4. ive gotten it down to S=2pi (x^3)/14 squareroot 1+(9x^4)/196 but now i am stuck

OpenStudy (anonymous):

I think i have to integrate again but i am not sure what to integrate

OpenStudy (anonymous):

Around which axis?

OpenStudy (anonymous):

x axis im sorry

OpenStudy (anonymous):

any ideas'

OpenStudy (anonymous):

Stay tuned

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[ 2 \pi f(x) \sqrt{f'(x)^2+1}=\frac{1}{7} \pi x^3 \sqrt{\frac{9 x^4}{196}+1}\\ \int \frac{1}{7} \pi x^3 \sqrt{\frac{9 x^4}{196}+1} \, dx=\frac{\pi \left(9 x^4+196\right)^{3/2}}{5292} \]

OpenStudy (anonymous):

and do i plug in 4 and 0 or do i need to find new limits of integration?

OpenStudy (anonymous):

\[ \int_0^4 \frac{1}{7} \pi x^3 \sqrt{\frac{9 x^4}{196}+1} \, dx=\frac{1132 \pi }{49} \]

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

No you plug in 4 and 0 and you substract.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

awesome!!!

OpenStudy (anonymous):

@eliassaab I am not getting the same answer. what equation do i plug those numbers into